OpenStudy (anonymous):

Derive y=(1/2)x^2 - In2x

OpenStudy (anonymous):

First, whats the derivative of (1/2) x^2?

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

yes! now whats the derivative of -ln2x

OpenStudy (anonymous):

-1/x?

OpenStudy (anonymous):

remember the chain rule ;)

OpenStudy (anonymous):

derivative of -ln2x = (-1/2x)(2) <--- chain rule the 2x

OpenStudy (anonymous):

now just put them together: x - 1/2x (2)

OpenStudy (anonymous):

Why is (-1/2x)(2) not the same as (-1/x)?

OpenStudy (anonymous):

oh i'm sorry you're absolutely right. My mind slipped for a second :(

OpenStudy (anonymous):

but you're getting the hang of it and even correcting me! so thats what matters right! :D

OpenStudy (anonymous):

I'm having a wrong answer with the stationary value then :/ Can you see if you get it?

OpenStudy (anonymous):

I thought my deriving was wrong... Must be the stationary value.

OpenStudy (anonymous):

what was the original question?

OpenStudy (anonymous):

The same as the last, the second question I asked there. Finding the stationary value + maxima and minima

OpenStudy (anonymous):

ok we know the derivative is x - 1/x right? so the stationary value is just when the derivative = 0, so we set x - 1/x = 0 0 = x - 1/x x = 1 my Calc 1 is a little rusty, I'll have to get back to you on how to find maximums and minimums :(

OpenStudy (anonymous):

Ok, so is the text book wrong? because I got 1 aswell. 1/2-In2 is what is says

OpenStudy (anonymous):

ohhh, because you take x=1 and plug it back into the original equation: y = (1/2) (x)^2 - ln(2x) y = (1/2) (1) - ln2 y = 1/2 - ln2 so the stationary POINT is (1, 1/2 - ln2) 1/2 - ln2 is the height of the point. Just looking at it, I would intuitively say its probably the minimum.

OpenStudy (anonymous):

Oh! Ok. That makes more sense. Thank you!!

OpenStudy (anonymous):

Yea no prob. I'm gonna head out now, so if you have anymore calculus questions ask me 2m :D

OpenStudy (anonymous):

Sure. Thanks again!